VISA Validation Suite USER MANUAL

Joo Thiam Goh Department of Ocean Engineering

Massachusetts Institute of Technology

Defence Science Organisation

20 Science Park Drive

Singapore 118230 Republic of Singapore

January 13, 1997

Contents

1 Testing the Installation 3 1.1 Installing the Benchmarks . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 Directory Tree . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 Running the scripts automatically . . . . . . . . . . . . . . . . . . . . . 4 1.4 Running the scripts manually . . . . . . . . . . . . . . . . . . . . . . . . 5

1 Testing the Installation

1.1 Installing the Benchmarks

The VISA distribution comes with a complete set of test problems and reference solutions. To install this set of scripts, follow the few simple steps :

• In the benchmarks directory, type

 make yesperl

or

 make noperl

depending on whether your particular system has PERL installed.

• Next you will need to compile the pltsplit program. Change to the ./src directory and type

 make

Then move the executable pltsplit to somewhere in your PATH .

• If you are using the perl-based scripts, you will also need to move ./bin/plpsub.pl to a location in your PATH .

1.2 Directory Tree

The scripts MUST be run from benchmarks . They assume that files are stored in the following manner :

./input              -> VISA input files
./logs               -> logs of each run
./refs/PE            -> reference solution for PE
./refs/BEM           -> reference solution for BEM
./refs/CORE          -> reference solution for CORE
./refs/COUPLE        -> reference solution for COUPLE
./refs/OASES         -> reference solution for OASES
./bin                -> All the relevant plotting scripts
./bin/YES_perl       -> Use these scripts if perl is present
./bin/NO_perl        -> Use these scripts if perl is not present
./tloss              -> Final combined PLT/PLP files
./tloss/default_plps -> PLP files for NO_perl scripts

    ./tloss/stubs        -> PLP template files for YES_perl scripts
    ./run                -> Results from VISA runs
    ./src                -> Source for pltsplit program

The VISA data files are located in the ./input directory. Since these represent the standard test set, DO NOT change the environmental variables in these files, as well as the number of field options. You may change the sampling parameters but if you did not install the perl-based scripts, then this might require that you change the combined PLP files in ./tloss after each run (see below).

1.3 Running the scripts automatically

To run a particular test simply call the appropriate script. For example to run Mike Collins Case D,

   do mikeExD

The final combined PLP/PLT files can be found in ./tloss/mikeExD.pl? . To run ALL the tests do

   doALL

These tests are run SEQUENTIALLY, i.e. one at a time to avoid starting multiple copies of VISA!! The doALL script takes an argument Q or q to indicate Quick runs, i.e. when invoked as

   doALL Q or doALL q

the script skip runs that involved more than 3 sectors.

Upon completion, the results of individual VISA runs are stored in ./run . The directory ./tloss holds the combined PLP/PLT files, i.e. they contain the VISA solutions as well as any other ’correct’ solutions for the problem. The plots are always ordered as OASES, BEM, VISA and CORE. Note that we may not have all the solutions from the various codes for each test case. Table 1 shows the available reference solutions for each of the different test case.

The above scripts execute the entire run, split and packaging for you. The next section describes the individual steps involved. Look at do mikeExD to get an idea of it.

Note that the basename for each test is used consistently for the input data file, scripts and combined PLP/PLT files. For example, do mikeExD , mikeExD.dat ,

./bin/mikeExD, ./tloss/mikeExD.pl*, ./logs/mikeExD.out, ./run/mikeExD.pl*

forms the complete set of files for running the Mike Collin Case D problem. Here, the basename is mikeExD . TABLE 1 contains a short description of the test cases.

1.4 Running the scripts manually

When VISA is done, you will need to split the PLT file, say X.plt into separate parts so that they can be merged with the reference solutions in ./refs/* into a single PLT file. The program that does this is pltsplit . The do * scripts calls pltsplit to do this for you automatically. The program pltsplit can be used in various manners.

• Interactively - simply call pltsplit and answer the questions.

• Via command line arguments - this is useful for automating the split in shell scripts (see do * ). The order of the input parameters is CRUCIAL.

• Type pltsplit -h to see the format of the command line inputs.

pltsplit splits BEM, VISA, SAFARI or CORE files into individual traces depending on the number of parameters (N,V,H etc), number of receiver depths and whether we have the forward, backward or total field components. It uses a very simple line counting algorithm and then split the file up based on the information you provided - robustness is questionable but so far it works. It then creates a series of files with unique filenames. For example, if you have N,V,H and 2 receivers with forward, backward and total field components, the file X.plt will be split into

X.rdo.par1.for.rx1.dat    X.rdo.par1.bac.rx1.dat
X.rdo.par2.for.rx1.dat    X.rdo.par2.bac.rx1.dat
X.rdo.par3.for.rx1.dat    X.rdo.par3.bac.rx1.dat
X.rdo.par1.for.rx2.dat    X.rdo.par1.bac.rx2.dat
X.rdo.par2.for.rx2.dat    X.rdo.par2.bac.rx2.dat
X.rdo.par3.for.rx2.dat    X.rdo.par3.bac.rx2.dat
X.rdo.par1.tot.rx1.dat
X.rdo.par2.tot.rx1.dat
X.rdo.par3.tot.rx1.dat
X.rdo.par1.tot.rx2.dat
X.rdo.par2.tot.rx2.dat
X.rdo.par3.tot.rx2.dat

The extensions for, bac, tot stands for forward, backward and total field components. The extensions rdo, gem, saf stands for VISA, CORE and SAFARI files.

1 TESTING THE INSTALLATION 6

Input file (.dat ) DESCRIPTION  REFERENCE
RSWT2a  Reverb. + Scat. Workshop Test case 2a COUPLE, CORE
RSWT2b  Reverb. + Scat. Workshop Test case 2b COUPLE, CORE
RSWT3a  Reverb. + Scat. Workshop Test case 3a COUPLE, CORE
RSWT3b  Reverb. + Scat. Workshop Test case 3b COUPLE, CORE
ASAwedgeE  ASA Elastic wedge problem  PE, CORE
torusF  Embedded Torus (FLUID)  BEM, CORE
torusE  Embedded Torus (ELASTIC)  BEM, CORE
gilbert  Gilbert-Evans Fluid  CORE
gilbertE  Gilbert-Evans Elastic  CORE
HorizForce  E/E Test - Horizontal Point force.  BEM, CORE
VertForce  E/E Test - Vertical Point force.  BEM, CORE
PntSrc  E/E Test - Point source  BEM, CORE
TranspHF  Transparent Elastic - Horizontal Force  BEM, CORE
TranspVF  Transparent Elastic - Vertical Force  BEM, CORE
mikeExD  Mike Collins Case D  BEM, CORE
mikeExE  Mike Collins Case E  BEM, CORE
30degLBeam  30-deg Llyod mirror beam F-S CORNER BEM, CORE
45degLBeam  45-deg Llyod mirror beam F-S CORNER BEM, CORE
SLeeHC  Single-layer High-contrast E/E test  BEM, CORE
SLeet  Single-layer Transparent E/E test  BEM, CORE
SLeed  Single-layer E/E test  BEM, CORE
SLef  Single-layer E/F test  BEM, CORE
SLfe  Single-layer F/E test  BEM, CORE
norda3a  Modified NORDA 3a - Isovelocity case  OASES, CORE
norda3a25c2  Modified NORDA 3a - c^{2} case  OASES
fnuw  Single-layer Fluid transparent problem  OASES, CORE
SACL circle F  SACLANT semi-circle ridge (Fluid)  COUPLE, CORE
SACL step F  SACLANT step discontinuity (Fluid)  COUPLE, CORE

Table 1: One-line description for the test cases BEM is treated as SAFARI. The extensions par1, par2, par3 refers to N, V and H. Finally, the extensions rx*, refer to the receiver depths, with the shallow receiver designated first.

WARNING : To properly use the scripts of course assume that you always run with the same set of parameters and receiver depths. Also NO INTEGRAND PLOTS are expected. Table 2 shows the expected OPTIONS and receiver depths for the various test cases.

After the PLT file has been split using pltsplit , they can then be combined using the scripts in ./bin

AGAIN this assumes that the test case has been run using certain wavenumber integration parameters that results in a PLT plot of certain size. If this has been changed, you will need to change the combined PLP file in ./tloss (unless you are using the PERL-based scripts). TABLE 2 shows the default PLT file size for the VISA runs. Note that the scripts, input data file and the combined PLP files all have the same basename.

We have included PERL-based scripts that automatically extract the PLP information from the VISA runs and merge them into the combined PLP files. This bypassed the above step of manually changing the combined PLP files whenever the PLT size is different. However it means that PERL must be installed on your system. If PERL is not available, simply install the alternative set of scripts in ./bin/NO perl . See the section on Installing the Benchmarks for more information.

These scripts call mplot by default and has a silent mode in which the combined PLT is created but mplot is not called. The silent mode (call scripts with argument S ) is used in the do * script to automatically create the combined PLT file in ./tloss after the run is completed. The default mode is useful for single runs. The scripts also saved the previous combined PLP/PLT files with an extension .old . This allows you to recover the old results if the new ones are worse! (Very useful)

NOTE : These scripts assume that tcsh is located in /usr/local/bin . (see the first line of each script). If you are unable to run the scripts, check to make sure that tcsh is where it should be. If not simply create a link to it (you will need to be root to do this) ,

ln -sf wherever_tcsh_is /usr/local/bin/tcsh

1 TESTING THE INSTALLATION 8

Input file (.dat ) OPTIONS  Receiver depths (m) # PLT points
RSWT2a  N T J B  45  600
RSWT2b  N T J B  45  600
RSWT3a  P N T J B  45  600
RSWT3b  P N T J B  45  602
ASAwedgeE  K T C J  30, 150  1032
torusF  P N T J B  50, 170  549
torusE  P N T J B  50, 170  549
gilbert  N T J B  50  750
gilbertE  K S T J B  80  750
HorizForce  N V H T J P B 3  70  503
VertForce  N V H T J P B 2  35  503
PntSrc  N V H T J P B  35  503
TranspHF  N V H T J P B 2  35  471
TranspVF  N V H T J P B 2  35  471
mikeExD  P N T J B  100, 300  500
mikeExE  P N T J B  100, 300  1250
30degLBeam  P N T J B  100, 300  267
45degLBeam  P N T J B  100, 300  179
SLeeHC  P N T J B  35  2009
SLeet  P N V H T J B  35  1883
SLeed  P N T J B  35  1831
SLef  P N T J B  35  503
SLfe  P N T J B  35  503
norda3a25  P N T J B  50, 110  1250
norda3a25c2  P N T J F  50, 110  658
fnuw  P N V H T J B  30  654
SACL circle F  P N T J B  50, 170  191
SACL step F  P N T J B  50, 170  801

Table 2: Default parameter set up for the various test cases